How do you solve x-4/4=9/8
Answer by stanbon(75887)

solve x-4/4=9/8
-------------------
Multiply both sides by 4 to get:
---
x-4 == 4(9/8)
x-4 = 9/2
---
x = (8/2)+(9/2)
x = 17/2
x = 8 1/2
===================
Cheers,
Stan H.